### v4

@@ Abstract
We study a fundamental problem in Computational Geometry, the planar two-center problem. In this problem, the input is a set $S$ of $n$ points in the plane and the goal is to find two smallest congruent disks whose union contains all points of $S$. A longstanding open problem has been to obtain an $O(n\log n)$-time algorithm for planar two-center, matching the $\Omega(n\log n)$ lower bound given by Eppstein [SODA’97]. Towards this, researchers have made a lot of efforts over decades. The previous best algorithm, given by Wang [SoCG’20], solves the problem in $O(n\log^{2}n)$ time. In this paper, we present an $O(n\log n)$-time (deterministic) algorithm for planar two-center, which completely resolves this open problem.
@@ 1 Introduction
Given a set $S$ of $n$ points in the plane, the planar $k$-center problem asks for $k$ smallest congruent disks in the plane whose union contains $S$. Planar $k$-center is NP-hard if $k$ is given as part of the input. The best known algorithm for the problem takes $n^{O(\sqrt{k})}$ time [19].
@@
In the special case where $k=1$, the problem can be solved in linear time [7, 12, 23, 24], and this result not only holds in the plane but also in any fixed dimension. When $k=2$, the problem becomes substantially more challenging. First of all, unlike the one-center case, the two-center problem requires $\Omega(n\log n)$ time to solve in the algebraic decision tree model, as shown by Eppstein [15] by a reduction from the max-gap problem. A longstanding open problem has been to prove a matching upper bound for the problem, i.e., obtain an $O(n\log n)$-time algorithm. Towards this, researchers have made a lot of efforts over decades. The first published result on planar two-center appeared in 1991 by Hershberger and Suri [18], who considered the decision version of the problem only and presented an $O(n^{2}\log n)$-time algorithm, and the algorithm was subsequently improved to $O(n^{2})$ time by Hershberger [17]. Using the decision algorithm in [17] and the parametric search technique [22], Agarwal and Sharir [2] presented an $O(n^{2}\log^{3}n)$-time algorithm for the (original) two-center problem. The running time was slightly improved in a sequence of work [14, 20, 21]. A major breakthrough was achieved by Sharir [27], who proposed an $O(n\log^{9}n)$-time algorithm. This is the first algorithm that solves the problem in near-linear time (in fact, even the first algorithm with subquadratic runtime). Eppstein [15] and Chan [6] further improved this result, obtaining an $O(n\log^{2}n)$-time randomized algorithm and an $O(n\log^{2}n\log\log n)$-time deterministic algorithm, respectively. Since then, no progress had been made for over two decades, until recently in SoCG 2020 Wang [29] gave an $O(n\log^{2}n)$-time deterministic algorithm [^4] Independently, Tan and Jiang [28] also claimed a simple $O(n\log^{2}n)$-time algorithm, but unfortunately that algorithm was later found to be incorrect [29]..
@@
Choi and Ahn [8] considered a special case of the two-center problem where the two disks in an optimal solution overlap and a point in their intersection is given in the input (we refer to it as the anchored case); they gave an $O(n\log n)$-time algorithm by improving the parametric search scheme in [29], where an algorithm of $O(n\log n\log\log n)$ time is provided for this case. In addition, Choi and Ahn’s technique also solves in $O(n\log n)$ time the convex case in which points of $S$ are in convex position; this again improves the $O(n\log n\log\log n)$-time result in [29]. Albeit these progress on special cases, prior to this work, Wang’s $O(n\log^{2}n)$-time (deterministic) algorithm [29] remains the best known result for the general planar two-center problem.
@@
In this paper, we resolve this longstanding open problem completely, by presenting an $O(n\log n)$-time deterministic algorithm for the planar two-center problem. Remarkably, in terms of running time, this is the first improvement since the $O(n\log^{2}n)$-time (randomized) algorithm of Eppstein [15] in 1997. Our result follows from a new decision algorithm for the problem which runs in $O(n)$ time after $O(n\log n)$-time preprocessing, summarized in Theorem 1. Along with known techniques in the previous work, such a decision algorithm immediately leads to an $O(n\log n)$-time algorithmn for the two-center problem, as elucidated in the proof of Corollary 2.
Theorem 1.
Let $S$ be a set of $n$ points in the plane. After a preprocessing step on $S$ in $O(n\log n)$ time, for any given $r\geq 0$, one can compute in $O(n)$ time two congruent disks of radius $r$ in the plane that together cover $S$, or decide the nonexistence of such two disks.
Corollary 2.
@@
If $\delta\leq\frac{3}{2}\cdot\mathsf{opt}$, then the algorithm of Choi and Ahn [8] can compute an optimal solution for $S$ in $O(n\log n)$ time (in this case it is possible to sample $O(1)$ points so that at least one point is in the intersection of the two optimal disks; thus the problem reduces to the anchored case).
For the case $\delta>\frac{3}{2}\cdot\mathsf{opt}$, Eppstein [15] showed that if one can decide whether $r\geq\mathsf{opt}$ for any given $r$ in $T(n)$ time after $T_{0}(n)$-time preprocessing on $S$, then one can compute an optimal solution for $S$ in $O(T(n)\cdot\log n+T_{0}(n))$ time by Cole’s parametric search [10]. The algorithm of Theorem 1 solves this decision problem with $T(n)=O(n)$ and $T_{0}(n)=O(n\log n)$. Therefore, we can compute an optimal solution for $S$ in $O(n\log n)$ time.
Combining the two cases finally gives us an $O(n\log n)$-time algorithm for the planar two-center problem. Note that since we do not know which case holds for $S$, we run the algorithms for both cases and return the better solution. ∎
To prove Theorem 1, which is the main goal of this paper, our algorithm is relatively simple and elegant. However, the correctness analysis of the algorithm is technical, entailing a nontrivial combination of several novel and intriguing insights into the problem, along with known observations in the literature. Some of the new insights might be useful for other related problems as well.
@@
A number of variants of the two-center problem have also attracted much attention in the literature, e.g., [1, 3, 4, 11, 16, 25, 26, 31]. For example, if the centers of the two disks are required to be in $S$, the problem is known as the discrete two-center problem. Agarwal, Sharir, and Welzl [3] solved the problem in $O(n^{4/3}\log^{5}n)$ time; the logarithmic factor in the runtime was slightly improved recently by Wang [30]. An outlier version of the problem was studied in [1]. Other variants include those involving obstacles [16, 25, 26], bichromatic version [4, 31], and kinetic version [11], etc. It should be noted that although the general planar $k$-center problem is NP-hard [19], Choi, Lee, and Ahn recently gave a polynomial-time algorithm for the case where the points are given in convex position [9].
@@
The rest of the paper is organized as follows. In Section 2, we define some basic notion that will be used throughout the paper. Section 3 introduces a new concept, called $r$-coverage, which serves an important role in our algorithm. To the best of our knowledge, we are not aware of any previous work that used this concept to tackle the two-center problem before. We prove several properties about the $r$-coverage, which may be interesting in their own right. Our algorithm for Theorem 1 is presented in Section 4, while its correctness is proved in Section 5. Finally, in Section 6, we conclude the paper and pose some open questions.
@@ 2 Preliminaries
The $r$-circular hull of a point set $Q$ in the plane with respect to a value $r>0$ is defined as the common intersection of all radius-$r$ disks that contain $Q$ [13, 18]. We use $\alpha_{r}(Q)$ to denote the $r$-circular hull of $Q$. See Figure 1(a). Note that $Q$ is $r$-coverable iff $\alpha_{r}(Q)$ exists.
Wang [29] considered the problem of dynamically maintaining the circular hull of a set $Q$ of points in the plane under monotone insertions, that is, each inserted point is always to the right of all points in the current $Q$. The following result is obtained in [29].
Lemma 3 (Theorem 6.7 in [29]).
@@
For any $r>0$, an $r$-solution for a planar two-center instance $S$ is a set of two radius-$r$ disks in the plane whose union covers $S$; a solution for $S$ refers to an $r$-solution with some $r\geq 0$. We denote by $\mathsf{opt}(S)$ the minimum $r$ such that $S$ has an $r$-solution. An $r$-solution $\{D,D^{\prime}\}$ for $S$ is tight if $\mathsf{dist}(D,D^{\prime})$ is minimized (among all $r$-solutions). An $r$-solution $\{D,D^{\prime}\}$ for $S$ is $p$-anchored for a point $p\in\mathbb{R}^{2}$ if $p\in D\cap D^{\prime}$. The anchored two-center problem takes as input a set $S$ of points in the plane and a point $p\in\mathbb{R}^{2}$, and asks for a $p$-anchored $r$-solution for $S$ with the minimum $r$. As discussed in Section 1, Choi and Ahn [8] presented an $O(n\log n)$-time algorithm for the anchored problem.
Lemma 4 (Choi and Ahn [8]).
@@ 3 $\bm{r}$-Coverage
In the following, we prove several properties about $\mathcal{CR}_{r}(Q)$ based on an interesting relationship with $\alpha_{r}(Q)$. More specifically, we will show that $\mathcal{CR}_{r}(Q)$ is convex and $\partial\mathcal{CR}_{r}(Q)$ consists of circular arcs of radii $r$ and $2r$ alternatively. Roughly speaking, these arcs correspond to the vertices and arcs of $\alpha_{r}(Q)$ in the following way: Each arc of radius $r$ of $\partial\mathcal{CR}_{r}(Q)$ is an “antipodal arc” of an arc of $\alpha_{r}(Q)$ and each arc of radius $2r$ of $\partial\mathcal{CR}_{r}(Q)$ has a vertex of $\alpha_{r}(Q)$ as its center (see Figure 1(b)); further, the cyclic order of the vertices and arcs of $\alpha_{r}(Q)$ on $\partial\alpha_{r}(Q)$ is consistent with the order of their corresponding arcs of $\mathcal{CR}_{r}(R)$ on $\partial\mathcal{CR}_{r}(R)$. As such, given $\alpha_{r}(Q)$, $\mathcal{CR}_{r}(Q)$ can be constructed in time linear in the number of vertices of $\alpha_{r}(Q)$. We formally prove these properties below.
@@
For contradiction, we assume that $q$ is not on $\partial\mathcal{CR}_{r}(Q)$. Then, $q$ must be in the interior of $\mathcal{CR}_{r}(Q)$, since $q$ lies in $\mathcal{CR}_{r}(Q)$ by the definition of $\mathcal{CR}_{r}(Q)$. Consider the ray $\rho$ from $p$ to $q$. Since $q$ is in the interior of $\mathcal{CR}_{r}(Q)$, $\rho$ contains a point $q^{\prime}\in\mathcal{CR}_{r}(Q)$ such that $q$ is in the interior of $\overline{pq^{\prime}}$. Since $q$ is the antipodal point of $p$ on $\partial D$ and the radius of $D$ is $r$, we have $|pq|=2r$. Hence, $|pq^{\prime}|>2r$. Since $q^{\prime}\in\mathcal{CR}_{r}(Q)$, there exists a radius-$r$ disk $D^{\prime}$ containing $Q$ and $q^{\prime}$. Since $D^{\prime}$ contains $Q$, $D^{\prime}$ contains $\alpha_{r}(Q)$ as well. As $p\in\alpha_{r}(Q)$, we obtain that $D^{\prime}$ contains both $p$ and $q^{\prime}$. But this is impossible as the radius of $D^{\prime}$ is $r$ while $|pq^{\prime}|>2r$. ∎
For any point $p\in\partial\alpha_{r}(Q)$, define $\tau(p)$ as the set of points $q$ satisfying conditions of Lemma 6, i.e., there is a radius-$r$ disk $D$ containing $Q$ and having both $p$ and $q$ on its boundary as antipodal points. By Lemma 6, $\tau(p)\subseteq\partial\mathcal{CR}_{r}(Q)$.
Let $e$ be an arc of $\alpha_{r}(Q)$ with two vertices $u$ and $v$. By slightly abusing the notation, define $\tau(e)=\cup_{p\in e}\tau(p)$. Let $D$ be the unique radius-$r$ disk whose boundary contains $e$. Note that $D$ contains $\alpha_{r}(Q)$ and thus contains $Q$ [13, 18, 29]. This property observes that $\tau(e)$ consists of the antipodal points of all points of $e$ on $\partial D$. We call $\tau(e)$ the antipodal arc of $e$. Since we have assumed that the distance of any two points of $Q$ is less than $2r$, $\tau(e)$ does not intersect $e$ [13, 18, 29]. For the reference purpose later, we have the following corollary, following directly from Lemma 6.
Corollary 7.
For an arc $e$ of $\alpha_{r}(Q)$, $\tau(e)$ is an arc lying on $\partial\mathcal{CR}_{r}(Q)$.
@@
Let $u$ be a vertex of $\alpha_{r}(Q)$. We now give a characterization of $\tau(u)$. Let $v$ and $w$ be the counterclockwise and clockwise neighboring vertices of $u$ on $\alpha_{r}(Q)$, respectively. See Figure 2. Let $e(u,v)$ be the arc of $\alpha_{r}(Q)$ connecting $u$ and $v$. Let $\hat{u}_{v}$ be the antipodal point of $u$ on the circle containing $e(u,v)$. Define $e(u,w)$ and $\hat{u}_{w}$ similarly. Observe that $\tau(u)$ is the arc from $\hat{u}_{v}$ clockwise to $\hat{u}_{w}$ on the radius-$(2r)$ circle centered at $u$. Also observe that the clockwise endpoint of $\tau(u)$, which is $\hat{u}_{w}$, is actually the counterclockwise endpoint of $\tau(e(u,w))$; similarly, the counterclockwise endpoint of $\tau(u)$, which is $\hat{u}_{v}$, is the clockwise endpoint of $\tau(e(u,v))$. Hence, if we consider the arcs $e$ and vertices $u$ of $\alpha_{r}(Q)$ in cyclic order on $\partial\alpha_{r}(Q)$, their corresponding arcs $\tau(\cdot)$ form a closed cycle, denoted by $\Lambda_{r}(Q)$. By Lemma 6, $\Lambda_{r}(Q)\subseteq\partial\mathcal{CR}_{r}(Q)$. In the following, we show that $\partial\mathcal{CR}_{r}(Q)\subseteq\Lambda_{r}(Q)$ (and thus $\partial\mathcal{CR}_{r}(Q)=\Lambda_{r}(Q)$). This is almost an immediate consequence of the following lemma.
Lemma 8.
@@
We first show that the antipodal point $p$ of $q$ on $\partial D$ lies on $\partial\alpha_{r}(Q)$. Assume to the contrary that this is not the case. Then, $p$ must lie outside of $\alpha_{r}(Q)$ since $\alpha_{r}(Q)\subseteq D$. Thus there is another disk $D^{\prime}$ of radius $r$ containing $Q$ but not containing $p$. As such, $Q\subseteq D\cap D^{\prime}$ and $p\notin D\cap D^{\prime}$. See Figure 3. Since $Q\subseteq D^{\prime}$, we have $D^{\prime}\subseteq\mathcal{CR}_{r}(Q)$. As $q\in\partial\mathcal{CR}_{r}(Q)$, $q$ cannot be in the interior of $D^{\prime}$. Further, since $p$ and $q$ are antipodal points on $\partial D$, it is possible to slightly change $D$ to a new position to obtain another radius-$r$ disk $D^{\prime\prime}$ such that $D^{\prime\prime}$ contains $D\cap D^{\prime}$ (and thus contains $Q$) and also contains $q$ in its interior (e.g., first slightly rotate $D$ around $q$ towards the interior of $D^{\prime}$, and then slightly translate the disk along the direction from the center toward $q$). Since $Q\subseteq D^{\prime\prime}$, it holds that $D^{\prime\prime}\subseteq\mathcal{CR}_{r}(Q)$. As $q$ is in the interior of $D^{\prime\prime}$, $q$ is also in the interior of $\mathcal{CR}_{r}(Q)$. But this incurs contradiction since $q$ lies on $\partial\mathcal{CR}_{r}(Q)$.
@@
We proceed to show that such a disk $D$ is unique. Assume there is another radius-$r$ disk $D^{\prime}$ with $Q\subseteq D^{\prime}$ and $q\in\partial D^{\prime}$. Then the antipodal point $p$ of $q$ on $\partial D$ lies outside of $D^{\prime}$. On the other hand, since $p$ lies on $\partial\alpha_{r}(Q)$ (which has been proved above), $p$ must be contained in $D^{\prime}$ because $D^{\prime}$ contains $Q$ (and thus contains $\alpha_{r}(Q)$). This leads to a contradiction. ∎
By Lemma 8, for any point $q\in\partial\mathcal{CR}_{r}(Q)$, $q$ must be in $\tau(p)$ for a point $p\in\partial\alpha_{r}(Q)$. As such, $q\in\Lambda_{r}(Q)$. This also proves that $\partial\mathcal{CR}_{r}(Q)\subseteq\Lambda_{r}(Q)$. We thus obtain $\partial\mathcal{CR}_{r}(Q)=\Lambda_{r}(Q)$. We summarize these in the following corollary (note that the arc $\tau(e)$ for each arc $e$ and $\tau(u)$ for each vertex $u$ of $\alpha_{r}(Q)$ can be computed in $O(1)$ time).
Corollary 9.
@@
The next lemma shows that $\mathcal{CR}_{r}(Q)$ is convex.
Lemma 10.

### v5

@@ Abstract
We study a fundamental problem in computational geometry, the planar two-center problem. In this problem, the input is a set $S$ of $n$ points in the plane and the goal is to find two smallest congruent disks whose union contains all points of $S$. A longstanding open problem has been to obtain an $O(n\log n)$-time algorithm for the planar two-center problem, matching the $\Omega(n\log n)$ lower bound given by Eppstein [SODA’97]. Towards this, researchers have made many efforts over decades. The previous best algorithm, given by Wang [SoCG’20], solves the problem in $O(n\log^{2}n)$ time. In this paper, we present an $O(n\log n)$-time (deterministic) algorithm for the planar two-center problem, which completely resolves this open problem.
@@ 1 Introduction
Given a set $S$ of $n$ points in the plane, the planar $k$-center problem asks for $k$ smallest congruent disks in the plane whose union contains $S$. Planar $k$-center is NP-hard if $k$ is given as part of the input. The best known algorithm for the problem takes $n^{O(\sqrt{k})}$ time [k-center].
@@
In the special case where $k=1$, the problem can be solved in linear time [ref:ChazelleOn96, ref:DyerOn86, ref:MegiddoLi83, Megiddo-linear], and this result not only holds in the plane but also in any fixed dimension. When $k=2$, the problem becomes substantially more challenging. First of all, unlike the one-center case, the two-center problem requires $\Omega(n\log n)$ time to solve in the algebraic decision tree model, as shown by Eppstein [eppstein1997faster] by a reduction from the max-gap problem. A longstanding open problem has been to prove a matching upper bound for the problem, i.e., obtain an $O(n\log n)$-time algorithm. Towards this, researchers have made many efforts over decades. The first published result on the planar two-center problem appeared in 1991 by Hershberger and Suri [HERSHBERGER1991431], who considered the decision version of the problem only and presented an $O(n^{2}\log n)$-time algorithm. This algorithm was subsequently improved to $O(n^{2})$ time by Hershberger [ref:HershbergerA93]. Using the decision algorithm in [ref:HershbergerA93] and the parametric search technique [parametric], Agarwal and Sharir [AgarwalPankajK1994Pglp] presented an $O(n^{2}\log^{3}n)$-time algorithm for the (original) two-center problem. The running time was slightly improved in a sequence of works [Eppstein-twocenter, JK-twocenter, KS-twocenter]. A major breakthrough was achieved by Sharir [Sharir-twocenter], who proposed an $O(n\log^{9}n)$-time algorithm. This is the first algorithm that solves the problem in near-linear time (in fact, even the first algorithm with subquadratic runtime). Eppstein [eppstein1997faster] and Chan [Chan-twocenter] further improved this result, obtaining an $O(n\log^{2}n)$-time randomized algorithm and an $O(n\log^{2}n\log\log n)$-time deterministic algorithm, respectively. Since then, no progress had been made for over two decades, until recently in SoCG 2020 Wang [wang2022planar] gave an $O(n\log^{2}n)$-time deterministic algorithm [^1] Independently, Tan and Jiang [ref:TanSi17] also claimed a simple $O(n\log^{2}n)$-time algorithm, but unfortunately that algorithm was later found to be incorrect [wang2022planar]..
@@
Choi and Ahn [choi2021efficient] considered a special case of the two-center problem where the two disks in an optimal solution overlap and a point in their intersection is given in the input (we refer to it as the anchored case); they gave an $O(n\log n)$-time algorithm by improving the parametric search scheme in [wang2022planar], where an algorithm of $O(n\log n\log\log n)$ time is provided for this case. In addition, Choi and Ahn’s technique also solves in $O(n\log n)$ time the convex case in which the points of $S$ are in a convex position; this again improves the $O(n\log n\log\log n)$-time result in [wang2022planar]. In spite of this progress on special cases, prior to this work, Wang’s $O(n\log^{2}n)$-time (deterministic) algorithm [wang2022planar] remains the best known result for the general planar two-center problem.
@@
In this paper, we resolve this longstanding open problem completely, by presenting an $O(n\log n)$-time deterministic algorithm for the planar two-center problem. Remarkably, in terms of running time, this is the first improvement since the $O(n\log^{2}n)$-time (randomized) algorithm of Eppstein [eppstein1997faster] in 1997, although Wang [wang2022planar] later presented a deterministic version with the same time complexity. Our result follows from a new decision algorithm for the problem which runs in $O(n)$ time after $O(n\log n)$-time preprocessing, summarized in Theorem 1.1 (unless otherwise stated, all time complexities in the following discussion are deterministic.) Along with known techniques in the previous works, such a decision algorithm immediately leads to an $O(n\log n)$-time algorithm for the two-center problem, as elucidated in the proof of Corollary 1.2.
Theorem 1.1.
Let $S$ be a set of $n$ points in the plane. After a preprocessing step on $S$ in $O(n\log n)$ time, for any given $r\geq 0$, one can compute in $O(n)$ time two congruent disks of radius $r$ in the plane that together covers $S$, or decide the nonexistence of two such disks.
Corollary 1.2.
@@
If $\delta\leq\frac{3}{2}\cdot\mathsf{opt}$, then the algorithm of Choi and Ahn [choi2021efficient] can compute an optimal solution for $S$ in $O(n\log n)$ time (in this case it is possible to sample $O(1)$ points in the plane (not necessarily from $S$) so that at least one point is in the intersection of the two optimal disks; thus the problem reduces to the anchored case).
For the case $\delta>\frac{3}{2}\cdot\mathsf{opt}$, Eppstein [eppstein1997faster] showed that if one can decide whether $r\geq\mathsf{opt}$ for any given $r$ in $T(n)$ time after $T_{0}(n)$-time preprocessing on $S$, then one can compute an optimal solution for $S$ in $O(T(n)\cdot\log n+T_{0}(n))$ time by Cole’s parametric search [ColeSl87]. The algorithm of Theorem 1.1 solves this decision problem with $T(n)=O(n)$ and $T_{0}(n)=O(n\log n)$. Therefore, we can compute an optimal solution for $S$ in $O(n\log n)$ time.
Combining the two cases finally gives us an $O(n\log n)$-time algorithm for the planar two-center problem. Note that since we do not know which case holds for $S$, we run the algorithms for both cases and return the better solution.
To prove Theorem 1.1, which is the main goal of this paper, we design an algorithm that is relatively simple. However, the correctness analysis of the algorithm is technical, entailing a nontrivial combination of several novel and intriguing insights into the problem, along with known observations in the literature.
@@
A number of variants of the two-center problem have also attracted much attention in the literature, e.g., [agarwal2008efficient, AgarwalPK1998TD2P, arkin2015bichromatic, de2013kinetic, halperin20022, OhBA19, OhCA18, wang2022improved]. For example, if the centers of the two disks are required to be in $S$, the problem is known as the discrete two-center problem. Agarwal, Sharir, and Welzl [AgarwalPK1998TD2P] solved the problem in $O(n^{4/3}\log^{5}n)$ time; the logarithmic factor in the runtime was slightly improved recently by Wang [ref:WangUn23]. An outlier version of the problem was studied in [agarwal2008efficient]. Other variants include those involving obstacles [halperin20022, OhBA19, OhCA18], the bichromatic version [arkin2015bichromatic, wang2022improved], the kinetic version [de2013kinetic], etc. It should be noted that although the general planar $k$-center problem is NP-hard [k-center], Choi, Lee, and Ahn recently gave a polynomial-time algorithm for the case where the points are given in convex position [ChoiLA23].
@@
The rest of the paper is organized as follows. In Section 2, we define some basic notions that will be used throughout the paper. Section 3 introduces a new concept, called $r$-coverage, which serves an important role in our algorithm. To the best of our knowledge, we are not aware of any previous work that used this concept to tackle the two-center problem before. We prove several properties about the $r$-coverage, which may be interesting in their own right. Our algorithm for Theorem 1.1 is presented in Section 4, while its correctness is proved in Section 5. Finally, in Section 6, we conclude the paper and pose some open questions.
@@ 2 Preliminaries
The $r$-circular hull of a point set $Q$ in the plane with respect to a value $r>0$ is defined as the common intersection of all radius-$r$ disks that contain $Q$ if such disks exist [ref:EdelsbrunnerOn83, HERSHBERGER1991431]. We use $\alpha_{r}(Q)$ to denote the $r$-circular hull of $Q$. See Figure 1(a). Note that $Q$ is $r$-coverable if and only if $\alpha_{r}(Q)$ exists and is not empty.
Wang [wang2022planar] considered the problem of dynamically tracking the circular hull of a set $Q$ of points in the plane under monotone insertions, that is, each inserted point is always to the right of all points in the current $Q$. The following result is obtained in [wang2022planar].
Lemma 2.1 (Theorem 6.7 in [wang2022planar]).
@@
More specifically, $\alpha_{r}(Q)$ is maintained in a linked list so that it can be output in time linear in its size [wang2022planar].
@@
For any $r>0$, an $r$-solution for a planar two-center instance $S$ is a set of two radius-$r$ disks in the plane whose union covers $S$; a solution for $S$ refers to an $r$-solution with some $r\geq 0$. We denote by $\mathsf{opt}(S)$ the minimum $r$ such that $S$ has an $r$-solution. An $r$-solution $\{D,D^{\prime}\}$ for $S$ is tight if $\mathsf{dist}(D,D^{\prime})$ is minimized (among all $r$-solutions). An $r$-solution $\{D,D^{\prime}\}$ for $S$ is $p$-anchored for a point $p\in\mathbb{R}^{2}$ if $p\in D\cap D^{\prime}$. The anchored two-center problem takes as input a set $S$ of points in the plane and a point $p\in\mathbb{R}^{2}$, and asks for a $p$-anchored $r$-solution for $S$ with the minimum $r$. As discussed in Section 1, Choi and Ahn [choi2021efficient] presented an $O(n\log n)$-time algorithm for the anchored problem.
Lemma 2.2 (Choi and Ahn [choi2021efficient]).
@@ 3 $\boldsymbol{r}$-Coverage
In the following, we prove several properties about $\mathcal{CR}_{r}(Q)$ based on an interesting relationship with $\alpha_{r}(Q)$. More specifically, we will show that $\mathcal{CR}_{r}(Q)$ is convex and $\partial\mathcal{CR}_{r}(Q)$ consists of circular arcs of radii $r$ and $2r$ alternately. Roughly speaking, these arcs correspond to the vertices and arcs of $\alpha_{r}(Q)$ in the following way: Each arc of radius $r$ of $\partial\mathcal{CR}_{r}(Q)$ is an “antipodal arc” of an arc of $\alpha_{r}(Q)$ and each arc of radius $2r$ of $\partial\mathcal{CR}_{r}(Q)$ has a vertex of $\alpha_{r}(Q)$ as its center (see Figure 1(b)); further, the cyclic order of the vertices and arcs of $\alpha_{r}(Q)$ on $\partial\alpha_{r}(Q)$ is consistent with the order of their corresponding arcs of $\mathcal{CR}_{r}(R)$ on $\partial\mathcal{CR}_{r}(R)$. As such, given $\alpha_{r}(Q)$, $\mathcal{CR}_{r}(Q)$ can be constructed in time linear in the number of vertices of $\alpha_{r}(Q)$. We formally prove these properties below.
@@
For contradiction, we assume that $q$ is not on $\partial\mathcal{CR}_{r}(Q)$. Then, $q$ must be in the interior of $\mathcal{CR}_{r}(Q)$, since $q$ lies in $\mathcal{CR}_{r}(Q)$ by the definition of $\mathcal{CR}_{r}(Q)$. Consider the ray $\rho$ from $p$ to $q$. Since $q$ is in the interior of $\mathcal{CR}_{r}(Q)$, $\rho$ contains a point $q^{\prime}\in\mathcal{CR}_{r}(Q)$ such that $q$ is in the interior of $\overline{pq^{\prime}}$. Since $q$ is the antipodal point of $p$ on $\partial D$ and the radius of $D$ is $r$, we have $|pq|=2r$. Hence, $|pq^{\prime}|>2r$. Since $q^{\prime}\in\mathcal{CR}_{r}(Q)$, there exists a radius-$r$ disk $D^{\prime}$ containing $Q$ and $q^{\prime}$. Since $D^{\prime}$ contains $Q$, $D^{\prime}$ contains $\alpha_{r}(Q)$ as well. As $p\in\alpha_{r}(Q)$, we obtain that $D^{\prime}$ contains both $p$ and $q^{\prime}$. But this is impossible as the radius of $D^{\prime}$ is $r$ while $|pq^{\prime}|>2r$.
For any point $p\in\partial\alpha_{r}(Q)$, define $\tau(p)$ as the set of points $q$ satisfying the conditions of Lemma 3.2, i.e., there is a radius-$r$ disk $D$ containing $Q$ and having both $p$ and $q$ on its boundary as antipodal points. By Lemma 3.2, $\tau(p)\subseteq\partial\mathcal{CR}_{r}(Q)$.
Let $e$ be an arc of $\alpha_{r}(Q)$ with two endpoints $u$ and $v$. By slightly abusing the notation, define $\tau(e)=\cup_{p\in e}\tau(p)$. Let $D$ be the unique radius-$r$ disk whose boundary contains $e$. Note that $D$ contains $\alpha_{r}(Q)$ and thus contains $Q$. Therefore, $\tau(e)$ consists of the antipodal points of all points of $e$ on $\partial D$. We call $\tau(e)$ the antipodal arc of $e$. Note that $\tau(e)$ and $e$ are congruent to each other (and thus they have the same length). Since we have assumed that the distance of any two points of $Q$ is less than $2r$, $\tau(e)$ does not intersect $e$. For later reference purposes, we have the following corollary, following directly from Lemma 3.2.
Corollary 3.4.
For an open arc $e$ of $\alpha_{r}(Q)$, $\tau(e)$ is an arc lying on $\partial\mathcal{CR}_{r}(Q)$.
@@
Let $u$ be a vertex of $\alpha_{r}(Q)$. We now give a characterization of $\tau(u)$. Let $v$ and $w$ be the counterclockwise and clockwise neighboring vertices of $u$ on $\alpha_{r}(Q)$, respectively. See Figure 2. Let $e(u,v)$ be the open arc of $\alpha_{r}(Q)$ connecting $u$ and $v$. Let $\hat{u}_{v}$ be the antipodal point of $u$ on the circle containing $e(u,v)$. Define $e(u,w)$ and $\hat{u}_{w}$ similarly. Observe that $\tau(u)$ is the arc from $\hat{u}_{v}$ clockwise to $\hat{u}_{w}$ on the radius-$(2r)$ circle centered at $u$. Also observe that the clockwise endpoint of $\tau(u)$, which is $\hat{u}_{w}$, is actually the counterclockwise endpoint of $\tau(e(u,w))$; similarly, the counterclockwise endpoint of $\tau(u)$, which is $\hat{u}_{v}$, is the clockwise endpoint of $\tau(e(u,v))$. Hence, if we consider the arcs $e$ and vertices $u$ of $\alpha_{r}(Q)$ in cyclic order on $\partial\alpha_{r}(Q)$, their corresponding arcs $\tau(\cdot)$ form a closed cycle, denoted by $\Lambda_{r}(Q)$. By Lemma 3.2, $\Lambda_{r}(Q)\subseteq\partial\mathcal{CR}_{r}(Q)$. In the following, we show that $\partial\mathcal{CR}_{r}(Q)\subseteq\Lambda_{r}(Q)$ (and thus $\partial\mathcal{CR}_{r}(Q)=\Lambda_{r}(Q)$). This is an almost immediate consequence of the following lemma.
Lemma 3.5.
@@
We first show that the antipodal point $p$ of $q$ on $\partial D$ lies on $\partial\alpha_{r}(Q)$. Assume to the contrary that this is not the case. Then, $p$ must lie outside of $\alpha_{r}(Q)$ since $\alpha_{r}(Q)\subseteq D$. Thus there is another disk $D^{\prime}$ of radius $r$ containing $Q$ but not containing $p$. As such, $Q\subseteq D\cap D^{\prime}$ and $p\notin D\cap D^{\prime}$. See Figure 3. Since $Q\subseteq D^{\prime}$, we have $D^{\prime}\subseteq\mathcal{CR}_{r}(Q)$. As $q\in\partial\mathcal{CR}_{r}(Q)$, $q$ cannot be in the interior of $D^{\prime}$. Further, since $p$ and $q$ are antipodal points on $\partial D$, it is possible to slightly change $D$ to a new position to obtain another radius-$r$ disk $D^{\prime\prime}$ such that $D^{\prime\prime}$ contains $D\cap D^{\prime}$ (and thus contains $Q$) and also contains $q$ in its interior (e.g., first slightly rotate $D$ around $q$ towards the interior of $D^{\prime}$, and then slightly translate the disk along the direction from the center toward $q$). Since $Q\subseteq D^{\prime\prime}$, it holds that $D^{\prime\prime}\subseteq\mathcal{CR}_{r}(Q)$. As $q$ is in the interior of $D^{\prime\prime}$, $q$ is also in the interior of $\mathcal{CR}_{r}(Q)$. But this leads to a contradiction since $q$ lies on $\partial\mathcal{CR}_{r}(Q)$.
@@
We proceed to show that such a disk $D$ is unique. Assume there is another radius-$r$ disk $D^{\prime}$ with $Q\subseteq D^{\prime}$ and $q\in\partial D^{\prime}$. Then the antipodal point $p$ of $q$ on $\partial D$ lies outside of $D^{\prime}$. On the other hand, since $p$ lies on $\partial\alpha_{r}(Q)$ (which has been proved above), $p$ must be contained in $D^{\prime}$ because $D^{\prime}$ contains $Q$ (and thus contains $\alpha_{r}(Q)$). This leads to a contradiction.
By Lemma 3.5, for any point $q\in\partial\mathcal{CR}_{r}(Q)$, $q$ must be in $\tau(p)$ for a point $p\in\partial\alpha_{r}(Q)$. As such, $q\in\Lambda_{r}(Q)$. This also proves that $\partial\mathcal{CR}_{r}(Q)\subseteq\Lambda_{r}(Q)$. We thus obtain $\partial\mathcal{CR}_{r}(Q)=\Lambda_{r}(Q)$. We summarize these in the following corollary (note that the arc $\tau(e)$ for each arc $e$ and $\tau(u)$ for each vertex $u$ of $\alpha_{r}(Q)$ can be computed in $O(1)$ time).
Corollary 3.7.
@@
Lemma 3.8.